confidence level and sample data to find confidence interval for estimating a population round your answer to the same number of decimal places as a sample

Question 2 2 pts Assume that a sample is used to estimate a population proportion p. Find the margin of error E that corresponds to the given statistics and confidence level. Round the margin of error to four decimal places. 90% confidence; n 390, x-146
Answer #1

p = x/n = 146/390 = 0.3744

Z for 90% confidence intercval = Z0.05 = 1.645

Margin of error = Z0.05 * sqrt(p * (1 - p) / n)

                         = 1.645 * sqrt(0.3744 * (1 - 0.3744) / 390)

                         = 0.0403